Determine whether the function f(x)=2x^2-x+2 as continuous at x=2
Answer by stanbon(75887)

Determine whether the function f(x)=2x^2-x+2 as continuous at x=2
1st: f(2) exists: f(2) = 2*2^2-2+2 = 8
2nd: show that the left limit as x approaches 2 is 8
3rd: show than the right limit as x appoaches 2 is 8
4th: conclude that f is continuous at x=2.
===================================================
Cheers,
Stan H.